# Learning Mean-Field Games through Mean-Field Actor-Critic Flow

### v1

@@ 2.1 Mean-field games
For fixed $\mu$, the problem (2.1)–(2.2) reduces to a classical stochastic control problem. Let $\alpha^{\mu,*}$ be the optimal control in this case, where the superscript $\mu$ emphasizes the dependence of $\alpha^{\mu,*}$ on the given flow of measure $\mu$. We denote the associated value function under this control by $V^{\mu,*}:=V^{\mu,\alpha^{\mu,*}}$. Then, by the dynamic programming principle, $V^{\mu,*}$ satisfies the HJB equation (cf. [55, Ch. 2-4])
@@ 4 The convergence analysis
$\mathcal{A}$ $:=\big\{\alpha:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{n}~|~~\alpha\text{ is twice differentiable in }x\in\mathbb{R}^{d},\,\,|\alpha(t,0)|\leq K,$
$\qquad\qquad|\alpha(t,x)-\alpha(s,x)|\leq K(1+|x|)|t-s|,\,\,|\nabla_{x}\alpha(t,x)|\leq K,~~|\nabla_{x}^{2}\alpha(t,x)|\leq K\big\},$
@@
for constants $c_{l},C_{l},c_{r},C_{r}>0$ depending only on $\sigma_{0}$, $d$, $K$ and $T$.
@@
$\mathcal{C}=\Big\{F(t,x)~\big|$ $\int_{\mathbb{R}^{d}}(1+|x|^{2})\,|F(t,x)|^{2}\rho(t,x)\,\mathrm{d}x\leq K\int_{\mathbb{R}^{d}}|F(t,x)|^{2}\rho(t,x)\,\mathrm{d}x,$ (4.2)
@@
$|\sigma(t,x,\mu)|,~|\nabla_{x}\sigma|\leq K,$
@@
Here, $\mu,\nu\in\mathcal{P}^{2}(\mathbb{R}^{d})$, and $\nabla_{x,\alpha}$ represents the gradient with respect to both $x$ and $\alpha$.
@@ Acknowledgment
The authors thank Alan Raydan for sharing the code associated with [4], which informed parts of our implementation. The authors thank Jiequn Han for making available the code from [28]; parts of this code were adapted for use in our experiments. The authors thank Daniel Lacker for discussion on MFGs.
@@
We provide technical lemmas and proofs used throughout the paper in the Appendix. $C$ and $c$ denote generic positive constants depending only on $d$, $K$, $T$, $\sigma_{0}$, $\lambda_{H}$, but independent of the speeds $\beta_{a}$, $\beta_{c}$, $\beta_{\mu}$. Their values may vary from line to line. $C$ is potentially large while $c$ is potentially small.
@@ A.1 Stochastic Grönwall’s inequalities
For $\alpha\in\mathcal{A}$, $\mu\in\mathcal{M}$, let $x_{t}:=X^{\mu,\alpha}_{t}$ be the state process under $(\mu,\alpha)$. There exists $C$ such that
$\mathbb{E}\big[|x_{t}|^{2}~|~x_{0}\big]\leq C(1+|x_{0}|^{2}).$ (A.1)
@@
If $\alpha=\alpha^{\prime}$, there exists $C$ such that
$\mathbb{E}\big[|x_{t}-x^{\prime}_{t}|^{2}~|~\mathcal{F}_{0}\big]\leq C(|x_{0}-x^{\prime}_{0}|^{2}+\mathcal{W}_{2}(\mu,\mu^{\prime})^{2}).$ (A.2)
If $\alpha=\alpha^{\prime}$ and $\mu=\mu^{\prime}$, there exists $C$ such that
$\mathbb{E}\Big[\left(1+|x_{t}|^{2}+|x^{\prime}_{t}|^{2}\right)\left|x_{t}-x^{\prime}_{t}\right|^{2}~\Big|~\mathcal{F}_{0}\Big]\leq C(1+|x_{0}|^{2}+|x_{0}^{\prime}|^{2})|x_{0}-x_{0}^{\prime}|^{2}.$ (A.3)
If $\alpha-\alpha^{\prime}\in\mathcal{C}$ and $x_{0}\overset{\mathrm{a.s.}}{=}x^{\prime}_{0}\sim\rho_{0}$, there exists $C$ such that
$\mathbb{E}\Big[\left(1+|x_{t}|^{2}+|x^{\prime}_{t}|^{2}\right)\left|x_{t}-x^{\prime}_{t}\right|^{2}\Big]\leq C\left(\mathcal{W}_{2}(\mu,\mu^{\prime})^{2}+\left\|\alpha-\alpha^{\prime}\right\|^{2}_{\mu,\alpha}\right).$ (A.4)
@@
$W_{2}(\rho^{\mu,\alpha}_{t},\rho^{\mu^{\prime},\alpha^{\prime}}_{t})^{2},\,\mathcal{W}_{2}(\rho^{\mu,\alpha},\rho^{\mu^{\prime},\alpha^{\prime}})^{2},\,d_{\beta}(\rho^{\mu,\alpha},\rho^{\mu^{\prime},\alpha^{\prime}})^{2}\leq C\left(\left\|\alpha-\alpha^{\prime}\right\|^{2}_{\mu,\alpha}+\mathcal{W}_{2}(\mu,\mu^{\prime})^{2}\right).$ (A.5)
@@
$\qquad\qquad\qquad\leq C\,\left(1+|x_{0}^{+}+x_{0}^{-}|^{2}+|x_{0}|^{2}\right)\left(|x_{0}^{+}+x_{0}^{-}-2x_{0}|^{2}+|x_{0}^{+}-x_{0}^{-}|^{4}\right).$ (A.8)
@@
$\mathbb{E}\Big[\big|(x^{2}_{t}-x^{1}_{t})-(x^{2\prime}_{t}-x^{1\prime}_{t})\big|^{2}~\big|~\mathcal{F}_{0}\Big]\leq C|x^{2}_{0}-x^{1}_{0}|^{2}\,\mathcal{W}_{2}(\mu,\mu^{\prime})^{2}.$ (A.9)
@@ A.3 Growth condition for the value function
$|V^{\mu,\alpha}(t,x)|,\,|\partial_{t}V^{\mu,\alpha}(t,x)|\leq C(1+|x|^{2}),$ (A.20)
$|\nabla_{x}V^{\mu,\alpha}(t,x)|\leq C(1+|x|),\quad|\nabla_{x}^{2}V^{\mu,\alpha}(t,x)|\leq C(1+|x|).$
@@ A.6 Moreau envelope
Let $R>1$ and $B_{R-1}$ be the closed ball in $\mathbb{R}^{d}$ of radius R1R-1 centered at origin. Assume that
$|V(t,x)|,|\partial_{t}V(t,x)|\leq C_{V}(1+|x|^{2}),\quad|\nabla_{x}V(t,x)|\leq C_{V}(1+|x|).$ (A.47)
The following conclusions hold:
@@
    For any $x\in B_{R-1}$ and $t\in[0,T]$, if $\iota\leq\frac{1}{4C_{V}R(2R^{2}+1)}\wedge\frac{1}{2(2C_{V}T+1)}$, then for any $(s,y)\in\textrm{Prox}_{\iota}[V](t,x)$ or $\textrm{Prox}^{\iota}[V](t,x)$,
@@
    $V(t,x)-V_{\iota}(t,x),V^{\iota}(t,x)-V(t,x)\in\left[0,8\iota\,C_{V}^{2}R^{2}(2R^{2}+1)\right].$ (A.49)
@@
    $V_{\iota}$ and $V^{\iota}$ satisfy a local Lipschitz condition: $\forall x,y\in B_{R-1}$, and $t,s\in[0,T]$,
@@
    The proximal operator satisfies the critical point equation: if $(s,y)\in\textrm{Prox}_{\iota}[V](t,x)$ or $\textrm{Prox}^{\iota}[V](t,x)$, then
    $\partial_{t}V_{\iota}(t,x)=\partial_{s}V(s,y),\quad\nabla_{x}V_{\iota}(t,x)=\nabla_{y}V(s,y),\quad\nabla_{x}^{2}V_{\iota}(t,x)\leq\nabla_{y}^{2}V(s,y),\quad\text{or}$ (A.52)
@@
We remark that the Moreau envelope is well studied, see [34] for detailed properties. Both $V_{\iota}$ and $V^{\iota}$ are well-defined whenever $\iota<1/(2C_{V})$. In (4) we assume $\iota<1/(2C_{V})$. By Alexandrov theorem [2], semiconvexity and semiconcavity imply that the second-order derivatives for $V_{\iota}$ and $V^{\iota}$ exist almost everywhere, so (A.52) holds almost everywhere.
@@
We show (1) first. Define $g_{\iota}(t,x):=V_{\iota}(t,x)-\frac{1}{2\iota}(t^{2}+|x|^{2})$. We show that $g_{\iota}(t,x)$ is concave. It is enough to verify that
$g_{\iota}(t+h,x+z)+g_{\iota}(t-h,x-z)\leq 2g_{\iota}(t,x),\quad\forall t,h,x,z,$
where $t-h,t+h]\subset[0,T]$. Assume $(s,y)\in\textrm{Prox}_{\iota}[V](t,x)$. Then,
@@
$=2V_{\iota}(t,x)-\tfrac{1}{\iota}(t^{2}+|x|^{2})=2g_{\iota}(t,x).$
This finishes (1).
We prove (2) next. Let $(s,y)\in\textrm{Prox}_{\iota}[V](t,x)$, then
$V(s,y)+\frac{1}{2\iota}(|t-s|^{2}+|x-y|^{2})\leq V(t,x).$
Therefore,
@@
which implies
$2R|t-s|+|x-y|\leq 4\iota\,C_{V}R(2R^{2}+1).$
This inequality, together with the range for $\iota$, further implies $|x-y|\leq 1$ and hence $|y|\leq R$. As a consequence,
$0\leq V(t,x)-V_{\iota}(t,x)=\frac{1}{2\iota}(|t-s|^{2}+|x-y|^{2})<2C_{V}R(2R|t-s|+|x-y|)\leq 8\iota C_{V}^{2}R^{2}(2R^{2}+1).$
This finishes proving (2).
We prove (3) next. By [34, Theorem 3.2], the superdifferential of $V_{\iota}(t,x)$ is the convex hull of $\frac{1}{\iota}[(t,x)-\textrm{Prox}_{\iota}[V](t,x)]$. Let $(s,y)$ lies in the convex hull of $\textrm{Prox}_{\iota}[V](t,x)$, then the estimates in (2) still hold (because quadratic function is convex), i.e.,
$|x-y|\leq 1,\quad|y|\leq R,\quad\frac{1}{2\iota}(|t-s|^{2}+|x-y|^{2})\leq 8\iota C_{V}^{2}R^{2}(2R^{2}+1).$
@@
We show (4) next. For $(s,y)\in\textrm{Prox}_{\iota}[V](t,x)$, using $|x-y|\leq 1$,
@@
Therefore, the map
$(\hat{t},\hat{x})\mapsto V_{\iota}(\hat{t},\hat{x})-V(\hat{t}-t+s,\hat{x}-x+y)$
attains its maximum at $(t,x)$. The first- and second-order optimal condition gives
$\partial_{t}V_{\iota}(t,x)=\partial_{s}V(s,y),\quad\nabla_{x}V_{\iota}(t,x)=\nabla_{y}V(s,y),\quad\nabla_{x}^{2}V_{\iota}(t,x)\leq\nabla_{y}^{2}V(s,y).$

### v2

@@ 2.1 Mean-field games
For fixed $\mu$, the problem (2.1)–(2.2) reduces to a classical stochastic control problem. Let $\alpha^{\mu,*}$ be the optimal control in this case, where the superscript $\mu$ emphasizes the dependence of $\alpha^{\mu,*}$ on the given flow of measure $\mu$. We denote the associated value function under this control by $V^{\mu,*}:=V^{\mu,\alpha^{\mu,*}}$. Then, by the dynamic programming principle, $V^{\mu,*}$ satisfies the HJB equation (cf. [56, Ch. 2-4])
@@ 4 The convergence analysis
$\mathcal{A}$ $:=\big\{\alpha:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{n}~|~~\alpha\text{ is twice differentiable in }x\in\mathbb{R}^{d},\,\,|\alpha(t,0)|\leq K,\,\,|\nabla_{x}\alpha(t,x)|\leq K,$
$|\nabla_{x}^{2}\alpha(t,x)|\leq K,\,\,|\alpha(t,x)-\alpha(s,x)|\leq K(1+|x|)\,\,|t-s|,\,\,|\nabla_{x}\alpha(t,x)-\nabla_{x}\alpha(s,x)|\leq K|t-s|,\big\},$
@@
for constants $c_{l},C_{l},c_{r},C_{r}>0$ depending only on $\sigma_{0}$, $d$, $K$ and $T$. In addition, we assume logarithmic Aronson bounds $\left|\nabla_{x}\log\rho^{\mu^{\tau},\alpha^{\tau}}(t,x)\right|\leq C(1+|x|)$ and $\left|\nabla_{x}^{2}\log\rho^{\mu^{\tau},\alpha^{\tau}}(t,x)\right|\leq C(1+|x|^{2})$, which will be used to prove a technical lemma in Appendix B.4. A similar bound was established in [51, Theorem B].
@@
$\mathcal{C}=\Big\{F(t,x)~\big|$ $\int_{\mathbb{R}^{d}}(1+|x|^{3})\,|F(t,x)|^{2}\rho(t,x)\,\mathrm{d}x\leq K\int_{\mathbb{R}^{d}}|F(t,x)|^{2}\rho(t,x)\,\mathrm{d}x,$ (4.2)
@@
$|\sigma(t,x,\mu)|,~|\nabla_{x}\sigma|,~|\nabla_{x}^{2}\sigma|\leq K,$
@@
Here, $\mu,\nu\in\mathcal{P}^{2}(\mathbb{R}^{d})$, and $\nabla_{x,\alpha}$ represents the gradient with respect to both $x$ and $\alpha$. In addition, we assume $\nabla_{x,\alpha}f$, $\nabla_{x,\alpha}^{2}f$, $\nabla_{x,\alpha}b$, $\nabla_{x,\alpha}^{2}b$, $\nabla_{x}\sigma$, $\nabla_{x}^{2}\sigma$, $g$, $\nabla_{x}g$ are all $K$-Lipschitz in $\mu$ with respect to $W_{2}(\cdot,\cdot)$.
@@ Acknowledgment
The authors thank Alan Raydan for sharing the code associated with [4], which informed parts of our implementation. The authors thank Jiequn Han for making available the code from [28]; parts of this code were adapted for use in our experiments. The authors thank Daniel Lacker for the discussion on MFGs.
@@
We provide technical lemmas and proofs used throughout the paper in the Appendix. Without specification, $C$ and $c$ denote generic positive constants depending only on $d$, $K$, $T$, $\sigma_{0}$, $\lambda_{H}$, but independent of the speeds $\beta_{a}$, $\beta_{c}$, $\beta_{\mu}$ and the selections of $\tau\geq 0,t\in[0,T],\ x\in\mathbb{R}^{d},\ \alpha\in\mathcal{A},\ \mu\in\mathcal{M}$. Their values may vary from line to line. $C$ is potentially large while $c$ is potentially small.
@@ A.1 Stochastic Grönwall’s inequalities
For $\alpha\in\mathcal{A}$, $\mu\in\mathcal{M}$, let $x_{t}:=X^{\mu,\alpha}_{t}$ be the state process under $(\mu,\alpha)$. Then
$\mathbb{E}\big[|x_{t}|^{2}~|~x_{0}\big]\leq C(1+|x_{0}|^{2}),\ \forall t\in[0,T].$ (A.1)
@@
If $\alpha=\alpha^{\prime}$,
$\mathbb{E}\big[|x_{t}-x^{\prime}_{t}|^{2}~|~\mathcal{F}_{0}\big]\leq C(|x_{0}-x^{\prime}_{0}|^{2}+\mathcal{W}_{2}(\mu,\mu^{\prime})^{2}),\ \forall t\in[0,T].$ (A.2)
If $\alpha=\alpha^{\prime}$ and $\mu=\mu^{\prime}$,
$\mathbb{E}\Big[\left(1+|x_{t}|^{2}+|x^{\prime}_{t}|^{2}\right)\left|x_{t}-x^{\prime}_{t}\right|^{2}~\Big|~\mathcal{F}_{0}\Big]\leq C(1+|x_{0}|^{2}+|x_{0}^{\prime}|^{2})|x_{0}-x_{0}^{\prime}|^{2},\ \forall t\in[0,T].$ (A.3)
If $\alpha-\alpha^{\prime}\in\mathcal{C}$ and $x_{0}\overset{\mathrm{a.s.}}{=}x^{\prime}_{0}\sim\rho_{0}$,
$\mathbb{E}\Big[\left(1+|x_{t}|^{2}+|x^{\prime}_{t}|^{2}\right)\left|x_{t}-x^{\prime}_{t}\right|^{2}\Big]\leq C\left(\mathcal{W}_{2}(\mu,\mu^{\prime})^{2}+\left\|\alpha-\alpha^{\prime}\right\|^{2}_{\mu,\alpha}\right),\ \forall t\in[0,T].$ (A.4)
@@
$W_{2}(\rho^{\mu,\alpha}_{t},\rho^{\mu^{\prime},\alpha^{\prime}}_{t})^{2},\,\mathcal{W}_{2}(\rho^{\mu,\alpha},\rho^{\mu^{\prime},\alpha^{\prime}})^{2},\,d_{\beta}(\rho^{\mu,\alpha},\rho^{\mu^{\prime},\alpha^{\prime}})^{2}\leq C\left(\left\|\alpha-\alpha^{\prime}\right\|^{2}_{\mu,\alpha}+\mathcal{W}_{2}(\mu,\mu^{\prime})^{2}\right),\ \forall t\in[0,T].$ (A.5)
@@
$\qquad\qquad\qquad\leq C\,\left(1+|x_{0}^{+}+x_{0}^{-}|^{2}+|x_{0}|^{2}\right)\left(|x_{0}^{+}+x_{0}^{-}-2x_{0}|^{2}+|x_{0}^{+}-x_{0}^{-}|^{4}\right),\ \forall t\in[0,T].$ (A.8)
@@
$\mathbb{E}\Big[\big|(x^{2}_{t}-x^{1}_{t})-(x^{2\prime}_{t}-x^{1\prime}_{t})\big|^{2}~\big|~\mathcal{F}_{0}\Big]\leq C|x^{2}_{0}-x^{1}_{0}|^{2}\,\mathcal{W}_{2}(\mu,\mu^{\prime})^{2},\ \forall t\in[0,T].$ (A.9)
@@ A.3 Growth condition for the value function
$|V^{\mu,\alpha}(t,x)|,\,|\partial_{t}V^{\mu,\alpha}(t,x)|\leq C(1+|x|^{2}),\quad|\nabla_{x}V^{\mu,\alpha}(t,x)|\leq C(1+|x|),$ (A.20)
$|\nabla_{x}^{2}V^{\mu,\alpha}(t,x)|\leq C(1+|x|),\ \forall(t,x)\in[0,T]\times\mathbb{R}^{d}.$
@@ A.6 Moreau envelope
Let $R>1$ and $B_{R-1}$ be the closed ball in $\mathbb{R}^{d}$ of radius R1R-1 centered at origin. Assume that there exists $C_{V}$ such that
$|V(t,x)|,|\partial_{t}V(t,x)|\leq C_{V}(1+|x|^{2}),\quad|\nabla_{x}V(t,x)|\leq C_{V}(1+|x|),\ \forall(t,x)\in[0,T]\times\mathbb{R}^{d}.$ (A.47)
The following conclusions hold under $\iota<\frac{1}{2C_{V}}$:
@@
    For any $x\in B_{R-1}$ and $t\in[0,T]$, if $\iota\leq\frac{1}{4C_{V}R(2R^{2}+1)}\wedge\frac{1}{4C_{V}(2T+1)}$, then for any $(s,y)\in\textrm{Prox}_{\iota}[V](t,x)$ or $\textrm{Prox}^{\iota}[V](t,x)$,
@@
    $V(t,x)-V_{\iota}(t,x),V^{\iota}(t,x)-V(t,x)\in\left[0,4\iota\,C_{V}^{2}R^{4}\right].$ (A.49)
@@
    $V_{\iota}$ and $V^{\iota}$ satisfy a local Lipschitz condition: for any $x,y\in B_{R-1}$ and $t,s\in[0,T]$,
@@
    The proximal operator satisfies the critical point equation: if $(s,y)\in\textrm{Prox}_{\iota}[V](t,x)$ resp. $\textrm{Prox}^{\iota}[V](t,x)$,
    $\partial_{t}V_{\iota}(t,x)=\partial_{s}V(s,y),\quad\nabla_{x}V_{\iota}(t,x)=\nabla_{y}V(s,y),\quad\nabla_{x}^{2}V_{\iota}(t,x)\leq\nabla_{y}^{2}V(s,y),\quad\text{resp.}$ (A.52)
@@
The Moreau envelope has been well studied in [34]. When $\iota<\frac{1}{2C_{V}}$, both envelopes $V_{\iota}$, $V^{\iota}$ are well-defined, and their associated proximal operators return non-empty sets. By Alexandrov theorem [2], either semiconvexity or semiconcavity implies the almost everywhere existence of the second-order derivatives of $V_{\iota}$ and $V^{\iota}$. As a result, (A.52) holds in the almost everywhere sense.
@@
We show (1) first. Define $g_{\iota}(t,x):=V_{\iota}(t,x)-\frac{1}{2\iota}(t^{2}+|x|^{2})$. We show that $g_{\iota}(t,x)$ is concave. It suffices to verify $g_{\iota}(t+h,x+z)+g_{\iota}(t-h,x-z)\leq 2g_{\iota}(t,x),\ \forall t,h,x,z$, where $t-h,t+h]\subset[0,T]$. For any $(s,y)\in\textrm{Prox}_{\iota}[V](t,x)$,
@@
$=2V_{\iota}(t,x)-\tfrac{1}{\iota}(t^{2}+|x|^{2})=2g_{\iota}(t,x),$
which concludes the proof.
Next, we prove (2). Let $(s,y)\in\textrm{Prox}_{\iota}[V](t,x)$, then $V(s,y)+\frac{1}{2\iota}(|t-s|^{2}+|x-y|^{2})\leq V(t,x).$ Therefore,
@@
which implies $2R|t-s|+|x-y|\leq 4\iota\,C_{V}R(2R^{2}+1).$ This inequality, together with the range for $\iota$, further implies $|x-y|\leq 1$ and hence $|y|\leq R$. As a consequence,
$0$ $\leq V(t,x)-V_{\iota}(t,x)=V(t,x)-V(s,y)-\frac{1}{2\iota}(|t-s|^{2}+|x-y|^{2})$
$\leq C_{V}\left[(1+R^{2})|t-s|+(1+R)|x-y|\right]-\frac{1}{2\iota}(|t-s|^{2}+|x-y|^{2})$
$\leq\frac{1}{2}\iota C_{V}^{2}[(1+R^{2})^{2}+(1+R)^{2}]\leq 4\iota C_{V}^{2}R^{4}.$
This concludes the proof of (2).
Next, we prove (3). By [34, Theorem 3.2], the superdifferential of $V_{\iota}(t,x)$ is the convex hull of $\frac{1}{\iota}[(t,x)-\textrm{Prox}_{\iota}[V](t,x)]$. Let $(s,y)$ lie in the convex hull of $\textrm{Prox}_{\iota}[V](t,x)$, then the estimates in (2) still hold, i.e.,
$|x-y|\leq 1,\quad|y|\leq R,\quad\frac{1}{2\iota}(|t-s|^{2}+|x-y|^{2})\leq 2C_{V}(2R^{2}|t-s|+R|x-y|)\leq 8\iota C_{V}^{2}R^{2}(2R^{2}+1).$
@@
Next, we show (4). For any $(s,y)\in\textrm{Prox}_{\iota}[V](t,x)$, using $|x-y|\leq 1$,
@@
Therefore, the mapping $(\hat{t},\hat{x})\mapsto V_{\iota}(\hat{t},\hat{x})-V(\hat{t}-t+s,\hat{x}-x+y)$ attains its maximum at $(t,x)$. The first- and second-order optimality condition provides
$\partial_{t}V_{\iota}(t,x)=\partial_{s}V(s,y),\quad\nabla_{x}V_{\iota}(t,x)=\nabla_{y}V(s,y),\quad\nabla_{x}^{2}V_{\iota}(t,x)\leq\nabla_{y}^{2}V(s,y),$
which concludes the proof. ∎